The coefficent of static friction between the floor of a truck and a box resting on it is 0.35. The truck is traveling at 81.3 km/hr. What is the least distance in which the truck can stop and ensure that the box does not slide?
Answer #1

acceleration till which the box friction can withstand

a = ug = 0.35* 9.8 = 3.43 m/s^2

now, using 3rd equation of motion

d = v^2/ (2 a)

d = (81.3* 5 / 18)^2 / ( 2* 3.43)

d = 74.345 m

=========

Comment before rate in case any doubt, will reply for sure.. goodluck